OpenStudy (anonymous):

simplify each expression 11-(d-2)+(-6)

OpenStudy (valpey):

Did you get the last one? Here you need to do things like distribute negative signs so it will be important to remember things like: -(a+b) = -a - b a+(-b) = a - b

OpenStudy (anonymous):

yes but this one is hard

OpenStudy (valpey):

Good luck! I'm sure you can do it. -(a+b) = -a - b a+(-b) = a - b

OpenStudy (anonymous):